OpenStudy (anonymous):

What is the area of a circle with a 13 cm diameter?

OpenStudy (anonymous):

132.73cm

OpenStudy (anonymous):

A=pi * r^2

OpenStudy (anonymous):

piR^2 13/2=6.5 6.5*6.5=42.25

OpenStudy (anonymous):

A≈132.73cm²

OpenStudy (anonymous):

42.25*3.14=132.665

OpenStudy (anonymous):

\[\pi \times 6.5^{2}\]\[6.5^{2}=42.25\]\[42.25\times \pi =132.665\]

OpenStudy (anonymous):

A≈132.73cm²

OpenStudy (anonymous):

shain show ur work where u got tht cuz they r saying differently

OpenStudy (anonymous):

Hippo how are you getting 132.73?

OpenStudy (anonymous):

hey plz don't strt a answering war its right

OpenStudy (anonymous):

look at ur equation again

OpenStudy (anonymous):

yah, PI*r^2

OpenStudy (anonymous):

look it up and see how u do it -_-

OpenStudy (anonymous):

Hippo got 132.73 because he probably just used Google calculator.

OpenStudy (anonymous):

lol i love it when u make tht face shain

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Lol, google says hippo is right. dk how, but it says so

OpenStudy (anonymous):

it doesn't matter how he got it it was right and yall need to stop depending on google

OpenStudy (anonymous):

im right right look atthe equation

OpenStudy (anonymous):

Although google is wrong. I did the same thing they did but still got 132.665

OpenStudy (anonymous):

No, Google is right, but when using Google, it uses the full pi instead of 3.14, therefore changing the answer slightly.

OpenStudy (anonymous):

Ahhh, I see

OpenStudy (anonymous):

well here is another one: Wht is the area of a circle with the radiuos of 53in

OpenStudy (anonymous):

53*53=2809 2809*3.14=8820.26